# GaussianOcc: Fully Self-supervised and Efficient 3D Occupancy Estimation with Gaussian Splatting

### v3

@@ 4.2 Implementation details
Network details: For U-Net architecture, we adapt New-CRFs [52] to predict the Gaussian attributes, which is based on the Swin Transformer [33]. The 6D pose network is the same as that used in SurroundDepth [44]. For the 2D-to-3D lifting, we follow the approach used in SimpleOcc [11]. In the depth estimation benchmark, we use the network proposed by SimpleOcc, where the final output size is 256×256×16. In our Gaussian splatting setting, we further upsample the final output to 512×512×32 for improved performance since we observe that a finer voxel grid leads to a finer rendered depth map, which requires ignorable computational cost. For occupancy estimation, we use the same network as OccNeRF [54] to ensure a fair comparison.
Training details: We propose a two-stage training process for fully self-supervised 3D occupancy estimation. In stage 1, we jointly train the depth estimation network and the 6D pose network, where we train the models for 8 epochs on the nuScenes dataset and 12 epochs on the DDAD dataset. In stage 2, we train the 3D occupancy network in a self-supervised manner with the 6D pose predicted from stage 1 rather than the ground truth pose used in OccNeRF [54]. We train the models for 12 epochs on both the nuScenes and DDAD datasets. The optimizer and learning rate adjustment strategy follow those used in SimpleOcc [11] and OccNeRF [54]. More implementation details (including code) are provided in the supplementary material.
@@ 4.3 Main results
3D occupancy estimation in nuScenes: As shown in Table 1, the proposed GaussianOcc achieves the best performance compared to other self-supervised methods. In particular, we do not require ground truth occupancy labels and ground truth 6D poses for training, thanks to the predicted 6D pose from the stage 1. In addition, we also achieve the good result on the RayIoU metric [31], with further details provided in the supplementary material.
3D occupancy estimation in DDAD: Thanks to our fully self-supervised learning setting, we are able to perform 3D occupancy estimation on the DDAD dataset [13]. To the best of our knowledge, our work is the first to introduce 3D occupancy estimation on this dataset, although we cannot provide quantitative results due to the lack of 3D occupancy label. We present visualization results in Figure 4. More results can be found in the supplementary material.
Depth estimation: We present a comparison of depth estimation results in Table 2 for both the nuScenes and DDAD datasets. In stage 1, GaussianOcc $\ddagger$ achieves top performance on the nuScenes dataset and delivers competitive results on the DDAD dataset. It is important to note that methods such as SurroundDepth [44] and SA-FSM [50] rely on third-party module for the sparse depth supervision. Additionally, SA-FSM is preprint work and has not released the code. R3D3 [38] is a temporal offline refinement method that requires multi-frame optimization, as discussed in [23]. In stage 2, which involves depth estimation from rendering, our method also achieves competitive results compared to those trained with ground truth poses. Besides, we observe that the rendered depth in stage two outperforms the depth results from stage 1 on the nuScenes dataset, whereas the opposite is true for the DDAD dataset. This discrepancy might be attributed to differences in perception range—80 meters in nuScenes versus 200 meters in DDAD. Moreover, an interesting phenomenon is that the semantic information is helpful for the depth estimation as indicated in GaussianOcc $\dagger$ whereas it worsens the result in OccNeRF $\dagger$. This shows the difference between the Gaussian splatting rendering and volume rendering.
@@ 4.4 Ablation Study
Scale-aware training: To demonstrate the effectiveness of the proposed scale-aware training using Gaussian Splatting, we conducted experiments comparing our method with existing approaches [44, 27] on Table 3. For a more fair comparison, we also implemented scale-aware training on [44, 27] that uses the same depth estimation network as ours [52], noted as $\checkmark$. We could reproduce the result on [27], but it is not good to training network with the sparse depth following [44]. Our proposed scale-aware training method is better than [44, 27]. Specifically, we observed that a naive implementation of Gaussian splatting without a masking strategy for cross-view rendering is ineffective because it would lead to a sub-optimal solution that the rendered image is still from the current view. Besides, to enhance performance, we introduced an erosion operation on the binary mask to purify it, excluding regions that may fall outside the overlap area. This step ensures better alignment during training. Finally, we refined the depth estimation by fixing the 6D pose network and disabling the cross-view loss. This refinement helps reduce artifacts at the edges of the overlap region. We present the cross-view splatting render result in Figure 5, where the rendering result in overlap regions is reasonable in the correct depth estimation from cross view that verifies our design.
6D pose learning and training strategy: Considering the GT pose from the sensor is imperfect as a lack of vertical movement [36]. We further evaluate the learned 6D pose quality by the self-supervised depth result at stage 2, where the experiments on Table 4 use the same network but with different pose constraints. Our result is competitive compared with GT pose, indicating that the predicted pose is of high quality. Furthermore, one stage training was ineffective in using the 6D pose from the jointly trained with 6D pose network because the cross-view loss in the 3D voxel space led to local optimization, which failed to generalize predictions to non-overlapping regions., as shown in Figure 6. However, this issue did not arise in the depth maps generated by the 2D decoder in stage 1, which is why we need the two-stage training process. Besides, our result is also better than [27], which is consistent the depth result on Table 3.
Volume rendering and Splatting rendering: In Table 5, we compare the performance of volume rendering (VR) and splatting rendering (SR) in the voxel space. For each 3D Gaussian, we set a uniform scale $s$ for all Gaussians, considering the well-arranged positions of vertices within the voxel grid. Since the scale $s$ is consistent across all three dimensions, we set the rotation $\mathbf{R}$ as the identity matrix. We first explored the optimal scale and found that a scale of 0.1 produced the best results. Notably, the performance of VR and SR was very close when using this scale. The render depth maps are shown in Figure 6. The depth map by the volume rendering is smoother thanks to dense sampling and the depth maps from splatting tendering have the graininess effect, especially at the small scale factor.
Pose and render types ablation study in occupancy task: In Table 6, we present an ablation study to evaluate the impact of using ground truth versus learned poses in the occupancy task, with training conducted using different rendering methods. The results show that splatting rendering achieves superior performance in both occupancy metrics (mIoU*) and depth metrics. One observation is that depth estimation results with the semantic learning in volume rendering are significantly worse than those in Table 2 whereas the proposed splatting rendering method maintains consistent performance, which suggests the splatting rendering in voxel grid contributes better 3D geometry.
Bonus of the fully self-supervised setting: The fully self-supervised setting of our method could be a general pretraining solution for supervised learning. After the self-supervised training on the DDAD and nuScenes datasets, we further finetune the model with the 3D occupancy label from Occ3D-nuScenes [41]. As shown in Table 8, experiments with self-supervised pretraining outperform the baseline. In particular, we find that pretraining on nuScenes is better than the DDAD dataset, which may own to the domain gap factors, such as differences in the scenarios (RGB images) and sensor configurations (camera extrinsics).
@@
Table 5: Comparison of the render result between the volume rendering (VR) [54] and splatting rendering (SR, Ours) on depth estimation task [14]. The coefficient $s$ means the scale of the 3D Gaussian. We use GT pose for this set of the ablation study.
@@
Table 7: Comparison of rendering efficiency between volume rendering (VR) [54] and splatting rendering (SR, Ours) on 3D occupancy estimation task [14]. The render time is calculated from surround 6 images. ”N/A” indicates out-of-memory errors running in NVIDIA A 100 (40 GB). Training time is averaged per epoch.
Method Pretraining setting mIoU
Baseline None 37.29
Self-supervised pretraining DDAD 37.40
nuScenes 38.45
Table 8: The study on SimpleOcc [11] with fully self-supervised pretraining. The baseline is directly training the model with 3D occupancy labels. The self-supervised pretraining is conducted on DDAD and nuScenes and then finetuned the model with 3D occupancy label. The number with bold typeface means the best.
Figure 5: Visualization of the depth map, splatting render image and depth map from cross view in stage 1.
@@ 5 Conclusion
In this paper, we introduce GaussianOcc, a fully self-supervised and efficient method for 3D occupancy estimation. Through the carefully designed cross-view splatting rendering, we can accurately learn the real scale in depth and the 6D pose, enabling effective self-supervised 3D occupancy learning. Additionally, the proposed Gaussian splatting in voxel grids outperforms volume rendering in 3D occupancy estimation while reducing computational cost. As a bonus, the fully self-supervised setting of the proposed method could serve as the general pretrain solution to enhance the 3D occupancy estimation performance. We validate our method on the nuScenes and DDAD datasets, demonstrating its strong generalization ability to surround view scenes.
6 More implementation details
@@
The detailed parameter setting in voxel grid splatting rendering. We chose a fixed scale for each grid vertex to ensure a well-arranged structure that accurately models the 3D space. If we use a learnable scale, it may lead to a situation where the scale is small but the opacity is large, which may not be captured in the rendered depth map and semantic map but could still affect the 3D occupancy result. Therefore, using a fixed scale is simple and sufficient for optimization, as demonstrated in the results presented in the main paper. Since the scale $s\in\mathbb{R}^{3}_{+}$ are identical for both three dimensions, we do not need to predict the rotation $r\in\mathbb{R}^{4}$ and set it with the identical matrix is sufficient. Similar to the OccNeRF [54], we render the 2D feature map for the semantic regression, while we leverage the 3D Gaussian splatting rendering and OccNeRF uses volume rendering.
@@
7 More experiment results
@@
More visualization. We provide more visualization for nuScenes dataset in Figure 7 and DDAD dataset in Figure 8. Please check the attached videos for sequence visualization.
More analysis on different supervision types. In Figure 9, we present visualizations of different supervision types. These visualizations highlight key differences in the results for the invisible regions (marked with red rectangles) and the rendered depth quality (marked with green rectangles).
@@
Additionally, the depth maps in the green-highlighted areas show that our Gaussian splatting rendering produces higher-quality depth predictions compared to volume rendering, consistent with the quantitative results in Table 6 of the main paper. Since Experiments (1) and (2) do not involve rendering-based training, they fail to produce reasonable depth predictions.
8 Limitation and future work
The proposed method achieves reasonable predictions in most scenes; however, we observe that some cases still present challenges, as shown in Figure 10. Specifically, in the DDAD dataset, incorrect predictions occur in the back camera in certain situations as marked with the red circle, where the drivable surface is mistakenly projected into the car due to extensive self-occlusion. Notably, this issue is absent in the nuScenes dataset, which has less self-occlusion. We believe that this problem could be mitigated with better 2D semantic maps for supervision, which warrants further investigation. The proposed method is for the surround view setting which is not suitable for the monocular images and the two-stage training procedure is a little bit time-consuming. Therefore, in the future, we will further investigate to apply the proposed method for one-stage training. Additionally, in stage 1, we leverage the spatial cross-view constraint through the proposed Gaussian splatting method. In the future, we aim to explore its potential benefits for temporal view synthesis as well.
@@ 8 Limitation and future work
Figure 8: The visualization of the render depth map and 3D occupancy prediction on DDAD dataset

### v4

@@ 4.2 Implementation details
Network details: For U-Net architecture, we adapt New-CRFs [52] to predict the Gaussian attributes, which is based on the Swin Transformer [33]. The 6D pose net is the same as that used in SurroundDepth [44]. For the 2D-to-3D lifting, we follow the approach used in SimpleOcc [11]. In the depth estimation benchmark, we use the network proposed by SimpleOcc, where the final output size is 256×256×16. In our Gaussian splatting setting, we further upsample the final output to 512×512×32 for improved performance since we observe that a finer voxel grid leads to a finer rendered depth map, which requires ignorable computational cost. For occupancy estimation, we use the same network as OccNeRF [54] to ensure a fair comparison.
Training details: We propose a two-stage training for fully self-supervised 3D occupancy estimation as indicated in Figure 2. In stage 1, we jointly train the depth estimation network and the 6D pose net, where we train the models for 8 epochs on the nuScenes and 12 epochs on the DDAD. In stage 2, we train the 3D occupancy network in a self-supervised manner with the 6D pose predicted from stage 1 rather than the ground truth pose used in OccNeRF [54]. We train the models for 12 epochs on both the nuScenes and DDAD. The optimizer and learning rate adjustment strategy follow those used in SimpleOcc [11] and OccNeRF [54].
@@ 4.3 Main results
3D occupancy estimation in nuScenes: In Table 1, the proposed GaussianOcc achieves the best performance compared to other self-supervised methods. In particular, we do not require ground truth occupancy labels and ground truth 6D ego pose for training, thanks to the predicted 6D pose from the stage 1. Note that RenderOcc [37] does not require the 3D occupancy label, but it is not a self-supervised method since it uses the ground truth depth map and semantic map for the 2D supervision; it could be regarded as a weakly-supervised method. In addition, we also achieve the best result on the RayIoU metric [31], with further details provided in the supplementary material.
3D occupancy estimation in DDAD: To the best of our knowledge, our work is the first to achieve 3D occupancy estimation on this dataset, thanks to our fully self-supervised learning setting. We present visualization results in Figure 4. As highlighted by the red rectangle, the sky region has a short-range depth value, but this does not appear in the rendered 3D occupancy estimation map thanks to the parameterized coordinate design of OccNeRF [54].
Depth estimation: We present a comparison of depth estimation results in Table 2 for both the nuScenes and DDAD. In stage 1, GaussianOcc $\ddagger$ achieves top performance on the nuScenes dataset and delivers competitive results on the DDAD. It is important to note that methods such as SurroundDepth [44] and SA-FSM [50] rely on third-party module for the sparse depth supervision. Additionally, SA-FSM is preprint work and has not released the code. R3D3 [38] is a temporal offline refinement method that requires multi-frame optimization, as discussed in [22]. In stage 2, which involves depth estimation from rendering, our method also achieves competitive results compared to those trained with ground truth poses. Besides, we observe that the rendered depth in stage two outperforms the depth results from stage 1 on the nuScenes, whereas the opposite is true for the DDAD. This discrepancy might be attributed to differences in perception range—80 meters in nuScenes versus 200 meters in DDAD. Moreover, an interesting phenomenon is that the semantic information is helpful for the depth estimation as indicated in GaussianOcc $\dagger$ whereas it worsens the result in OccNeRF $\dagger$. This phenomenon can be attributed to the biased sampling strategy of OccNeRF, where only 25% of the sample points are used for faster semantic map rendering compared to depth map rendering. In contrast, our proposed Gaussian splatting method, which renders directly from the voxel vertices, eliminates this issue.
Figure 5: The comparison of the depth map and its synthesis overlap image with (1) direct bilinear interpolation cross-view synthesis [44] and (2) our cross-view Gaussian splatting synthesis.
@@ 4.4 Ablation Study
Scale-aware training in Stage 1: We compare our method with existing approaches [44, 27] on Table 3 to demonstrate the effectiveness of the proposed scale-aware training using Gaussian Splatting. For a more fair comparison, we also implemented scale-aware training on [44, 27] that uses the same depth estimation network as ours [52], noted as $\checkmark$. We could reproduce the result on [27], but it is not good to training network with the sparse depth following [44]. Our proposed scale-aware training method is better than [44, 27]. Specifically, we observed that a naive implementation of Gaussian splatting without a mask-out strategy for cross-view rendering is ineffective because it would lead to a sub-optimal solution that the rendered image is still from the current view. Besides, to enhance performance, we introduced an erosion operation on the binary mask to purify it, excluding regions that may fall outside the overlap area. This step ensures better alignment during training. Finally, we refined the depth estimation by fixing the 6D pose net and disabling the cross-view loss. This refinement helps reduce artifacts at the edges of the overlap region. Visualization: In Figure 5, we can see that the depth map is easily be suboptimal depth estimation in near overlap region with the training of (1) direct cross-view interpolation constrain in [44]. Specifically, though the region with yellow rectangle has the incorrect large depth value, the synthesis region still presents a reasonable result that would lead to sub-optimal training. Thanks to our cross-view Gaussian splatting design analysis above, our method shown in (2) does not have the sub-optimal situation.
@@
SR ($s$ = learnable) 0.212 3.248 7.112 0.771
Table 5: Comparison of the render result between the volume rendering (VR) [54] and splatting rendering (SR, Ours) on depth estimation task [3]. The coefficient $s$ means the scale of the 3D Gaussian and learnable means the scale is learnable by the Sigmoid function. We use GT pose for this set of the ablation study.
@@
Table 7: Comparison of rendering efficiency between volume rendering (VR) [54] and splatting rendering (SR, Ours) on 3D occupancy estimation task [3]. The render time is calculated from surround 6 images. "N/A" indicates out-of-memory errors running in NVIDIA A 100 (40 GB). Training time is averaged per epoch.
6D pose learning and training strategy: Considering the 6D ego pose from the sensor is imperfect as a lack of vertical movement [36]. We further evaluate the learned 6D pose quality by self-supervised depth result at stage 2, where the experiments on Table 4 use the same network but with different pose constraints. Our result is competitive compared with GT pose, indicating that the predicted pose is of high quality. Furthermore, one stage training was ineffective in using the 6D pose from the jointly trained with 6D pose net because the cross-view loss in the 3D voxel space led to local optimization, which failed to generalize predictions to non-overlapping regions, as shown in Figure 6. However, this issue did not occur in the depth maps produced by the 2D decoder in Stage 1, highlighting the necessity of the two-stage training. Besides, our result is also better than [27], which is consistent the depth result on Table 3.
Volume rendering and Splatting rendering: In Table 5, we compare the performance of volume rendering (VR) and splatting rendering (SR) in the voxel space. For each 3D Gaussian, we first set a uniform scale $s$ for all Gaussians, considering the well-arranged positions of vertices within the voxel grid. Since the scale $s$ is consistent across all three dimensions, we set the rotation $\mathbf{R}$ as the identity matrix. We found that a scale of 0.1 produced the best results, closely matching the VR at this scale. Besides, by applying a learnable scale through the output of a Sigmoid function and clamping the maximum scale to 0.12, we achieved the highest performance. The render depth maps are shown in Figure 6. The depth map by the volume rendering is smoother thanks to dense sampling and the depth maps from splatting rendering have the graininess effect, especially at the small scale factor (0.05).
Pose and render types ablation study in occupancy task: In Table 6, we present an ablation study to evaluate the impact of using ground truth versus learned poses in the occupancy task, with training conducted using different rendering methods. The results show that splatting rendering achieves superior performance in both occupancy metrics (mIoU*) and depth metrics. One observation is that depth estimation results with the semantic learning in volume rendering are significantly worse than those in Table 2 whereas our proposed splatting rendering method maintains consistent performance, which suggests the splatting rendering in voxel grid contributes better 3D geometry.
@@ 5 Conclusion
In this paper, we introduce GaussianOcc, a fully self-supervised and efficient method for 3D occupancy estimation. Through the carefully designed cross-view splatting rendering, we can accurately learn the real scale in depth and the 6D pose, enabling effective self-supervised 3D occupancy learning. Additionally, the proposed Gaussian splatting in voxel grids outperforms volume rendering in 3D occupancy estimation while reducing computational cost.
Acknowledgments This work was supported in part by the JSPS, KAKENHI under Grant Number 22H03609, JST, FOREST under Grant Number JPMJFR206S. Wanshui Gan was also supported by RIKEN JRA Program. Big thanks for Xiaoyu Dong for helping proofreading the manuscript.
Appendix
Appendix A More implementation details
@@
The detailed parameter setting in voxel grid splatting rendering for semantic rendering. For semantic rendering, we chose a fixed scale for each grid vertex to ensure a well-arranged structure that accurately models the 3D space. If we use a learnable scale, it may lead to a situation where the scale is small but the opacity is large, which may not be captured in the rendered depth map and semantic map but could still affect the 3D occupancy result. Therefore, using a fixed scale is simple and sufficient for optimization, as demonstrated in the results presented in the main paper (Table 5) that the performance is close to the learnable scale. Since the scale $s\in\mathbb{R}^{3}_{+}$ are identical for both three dimensions, we do not need to predict the rotation $r\in\mathbb{R}^{4}$ and set it with the identical matrix is sufficient. Similar to the OccNeRF [54], we render the 2D feature map for the semantic regression, while we leverage the 3D Gaussian splatting rendering and OccNeRF uses volume rendering.
@@
Figure 7: One-stage training analysis. This is the visualization of the overlap mask and the rendered depth map with one-stage training.
Appendix B More experiment results and analysis
Why we need two-stage training. We made extensive efforts to develop one-stage training that directly applies the cross-view loss to the rendered depth map but were unsuccessful. As suggested in Figure 7, the cross-view supervision signals are effective only in overlapping regions. The rendered depth map learned from 3D CNN has lower generalization ability in non-overlap regions compared with the decoder depth learned by the 2D CNN, which led to local minima.
@@ Appendix B More experiment results and analysis
More visualization. We provide more visualization for nuScenes dataset in Figure 8. Please check the videos for sequence visualization in https://github.com/GANWANSHUI/GaussianOcc.git.
More analysis on 3D occupancy and depth map result on different supervision types.
3D occupancy analysis: In Figure 9, we present visualizations of different supervision types. These visualizations highlight key differences in the results for the invisible regions (marked with red rectangles) and the rendered depth quality (marked with green rectangles).
@@
Render depth map analysis: In Tables 2 and 6 of main paper, we observe an interesting phenomenon that the semantic information is helpful for the depth estimation with our GaussianOcc whereas it worsens the result in OccNeRF. In Figure 9, we visualize the depth map and highlight with green rectangles that our Gaussian splatting rendering produces higher-quality depth predictions compared to volume rendering. This should be concluded to the biased sampling strategy of OccNeRF, where only 25% of the sample points are used for faster semantic map rendering compared to depth map rendering. Here is the piece of the code in OccNeRF [54]. In contrast, our proposed Gaussian splatting method, which renders directly from the voxel vertices, eliminates this issue. At last, since Experiments (1) and (2) do not involve rendering-based training, they fail to produce reasonable depth predictions.
Render resolution Render time (s) with different voxel resolutions (Gaussians number)
$180\times 320$ $16\times 200\times 200$ $24\times 300\times 300$ $32\times 512\times 512$
VR $\approx$ 0.50 $\approx$ 0.85 $\approx$ 1.52
SR $\approx$ 0.06 $\approx$ 0.17 $\approx$ 0.44
Table 8: Comparison of rendering efficiency under different Gaussians number between volume rendering (VR) [54] and splatting rendering (SR, Ours).
Gaussians number and its related render time: (1) In stage 1, Gaussians number depends on the depth map resolution from the 2D decoder, where each pixel is a Gaussian primitive after unprojection. We use the depth map resolution in $224\times 352$, resulting in 78,848 Gaussian primitives in one image. In stage 2, Gaussians number depends on the voxel resolution, where each voxel grid is a Gaussian primitive. In Table 7 of the main paper, we follow the voxel resolution the same as OccNeRF [52] in $24\times 300\times 300$, resulting in 2,160,000 Gaussian primitives. (2) We revealed the rendering time under different render image resolutions compared with volume rendering in Table 7 of the main paper. We conducted the extra experiment for render time comparison under the same render image resolution ($180\times 320$) but with different Gaussians number (voxel resolutions) as shown in Table 8. From Table 7 of the main paper and Table 8, we observe: (1) The render time of splatting rendering (SR) is mainly affected by the Gaussians number, not the render image resolution. (2) SR is 3–8 times faster than volume rendering (VR) across different voxel settings.
Bonus of the fully self-supervised setting: The fully self-supervised setting of our method could be a general pretraining solution for supervised learning. After the self-supervised training on the DDAD and nuScenes datasets, we further finetune the model with the 3D occupancy label from Occ3D [41]. As shown in Table 11, experiments with self-supervised pretraining outperform the baseline. In particular, we find that pretraining on nuScenes is better than the DDAD dataset, which may own to the domain gap factors, such as differences in the scenarios (RGB images) and sensor configurations (camera extrinsics).
Appendix C Limitation and future work
The proposed method achieves reasonable predictions in most scenes; however, we observe that some cases still present challenges, as shown in Figure 10. Specifically, in the DDAD dataset, incorrect predictions occur in the back camera in certain situations as marked with the red circle, where the drivable surface is mistakenly projected into the car due to extensive self-occlusion. Notably, this issue is absent in the nuScenes dataset, which has less self-occlusion. We believe that this problem could be mitigated with better 2D semantic maps for supervision, which warrants further investigation. The proposed method is for the surround view setting which is not suitable for the monocular images. Additionally, in stage 1, we leverage the spatial cross-view constraint for scale-aware training through the proposed Gaussian splatting method. In the future, we aim to explore its potential benefits for temporal view synthesis as well.